If I am 35 years old and my son is 5 and how many years will I be exactly three times older than my son
Found 2 solutions by FrankM, macston:
Answer by FrankM(1040)

Some years from now is X years away. You will be 35+X and the boy, 5+X.
35+X = 3(5+X)
35+X = 15+3X
20 = 2X
X = 10
In 10 years you will be 45 and the young man, 15

Answer by macston(5194)

.
Y=your age=35; S=son's age=5
x=number of years until you are 3 times older
.
Y+x=3(S+x)
35+x=3(5+x)
35+x=15+3x
20=2x
10=x
.
ANSWER: You will be three times your son;s age in ten years.
(You will be 45 and he will be 15)
.